OpenStudy (anonymous):

If y=e^x^2 ln(x^2), find dy/dx

zepdrix (zepdrix):

\[\huge y=e^{x^2}\cdot \ln(x^2)\]That's our y? :D Oh boy.

zepdrix (zepdrix):

Sooooooo, product rule yes? We have 2 separate thingies with x in them.\[\large \frac{dy}{dx}=\frac{d}{dx}\left[e^{x^2}\right]\ln(x^2)+e^{x^2}\frac{dy}{dx}\left[\ln(x^2)\right]\]Understand the setup? :)

OpenStudy (anonymous):

not really

zepdrix (zepdrix):

I prefer to use the PRIME notation, it's a little easier to read, let's see if that makes a difference. Here is the definition of product rule,\[\huge (uv)'=u'v+uv'\]

zepdrix (zepdrix):

\[\large \left(e^{x^2}\cdot \ln(x^2)\right)'=\left(e^{x^2}\right)'\cdot \ln(x^2)+e^{x^2}\cdot\left(\ln(x^2)\right)'\] The primes are where we need to take a derivative.

zepdrix (zepdrix):

It's a bit of a tricky problem! :O If you can understand the setup then it's not too bad after that :D

OpenStudy (anonymous):

so how do i take the derivative of e^x^2 and ln(x^2)

zepdrix (zepdrix):

Remember that with an exponential of base e, the exponential won't change, but we'll apply the chain rule, multiplying by the derivative of the exponent that's up top. \[\huge (e^{x^2})'=e^{x^2}(x^2)'\] Giving us,\[\huge e^{x^2}(2x)\]Understand that one? :O

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Remember the derivative of natural log? :)\[\large (\ln (x)\;)'=\frac{1}{x}\]It's just important to remember that WHATEVER is inside of the log, is what you stuff into the denominator.\[\large (\ln(x^2)\;)'=\frac{1}{x^2}(x^2)'\]Then of course we have to apply the chain rule again :D\[\large =\frac{1}{x^2}(2x)\qquad =\qquad \frac{2}{x}\]